Status Report

NASA Space Station On-Orbit Status 4 Dec 2003

By SpaceRef Editor
December 4, 2003
Filed under , , ,
NASA Space Station On-Orbit Status 4 Dec 2003
iss

All ISS systems continue to function nominally, except as noted previously or below.  >>>Today five years ago, STS-88/Endeavour, commanded by Bob Cabana, launched the second ISS element, Node “Unity” with PMAs 1 & 2, on ISS Mission 2A (12/4/98).  Also today it is one solid year that the ISS community has relied entirely on our Russian partner for launch support.<<<

FE Alexander Kaleri conducted the first experiment session with the Russian TEKh-20 Plasma Crystal-3 (PK-3) payload.  [Right after wake-up, Sasha activated the turbopump in the Service Module’s Transfer Compartment (SM PkhO), tagged up with ground specialists and started the evacuation of the vacuum chamber (ZB) in the SM Work Compartment (RO), then monitored PK-3 operations.  At Experiment Start + 13 min., live TV downlink via S-band provided coverage for ~11 min.  The turbopump was to be deactivated tonight at ~4:25pm EST.  The experiment is performed on plasma, i.e., fine particles charged and excited by HF (high frequency) radio power inside the evacuated work chamber.  Main objective is to study dust plasma crystallization processes at a specified power of HF discharge, pressure, and a varied number of particles with subsequent reduction of HF discharge power, then to observe melting of the structures formed earlier.  The experiment is conducted in automated mode.]

CDR Michael Foale started early on a planned reloading of two SSC (station support computer) laptops, SSC6 and SSC2, which was necessary to allow IWIS software to continually operate during today’s EMU battery discharge ops, which also required an SSC. [On SSC2, Mike found its CD drive to be missing a bearing (both SSCs are IBM 760XD ThinkPads, not NGL/Next Generation Laptop IBM A31p TPs).  Using the CD drive from SSC3 instead, the computer was successfully reloaded.]

The CDR terminated the 24-hour charging process for EMU batteries #2032 and #2033 in the Airlock (A/L)’s battery stowage assembly (BSA), then initiating their discharge after the required one-hour rest period, during which the units remained in the BSA.  The standalone SSC7 laptop in the A/L needed for the procedure was first reconfigured for the discharge, then restored to its original settings. [The periodic (50-day) charge/discharge cycle is a common maintenance requirement to restore the batteries’ capacity, prolonging their useful life.]

After readying a number of empty EDV water containers, Kaleri filled them with water from the Progress-248/12P “Rodnik” tank (BV2) until it was drained.  The water can be used for any needs, including drinking.

Sasha completed the daily SOZh life support systems maintenance (including toilet facility, food containers, water containers and solid waste containers).  He also prepared the daily IMS (inventory management system) update file.

At 9:45am EST, both crewmembers participated in an interactive televised PAO interview (Ku-band and S-band) with England’s Sir David Frost for the British Broadcasting Company (BBC).  The interview, which went out live on NASA TV, was recorded for later playback on the high-profile BBC TV program “Breakfast with Sir David Frost”.

Mike Foale received thanks from POIC (Payload Operations & Integration Center, Huntsville) for successfully conducting yesterday’s FOOT experiment (Foot/Ground Reaction Forces during Space Flight).  [Status on the collected data will be provided after their downlink and analysis.  From video coverage, it appears that further alteration of the LEMS pants for proper electrode placement on a leg muscle may be required before next data collection.]

The hard drive (#6054) of an A/L PCS (portable computer system) was declared failed yesterday when it could not be recovered.  [There are no properly configured spare PCS drives on board, but there are three functioning PCS laptops (Flight Rule minimum is two).  Impact: loss of redundancy for crew insight, incl. Caution & Warning, and command interface to the U.S. segment.  Spare hard drives will be suitably configured (“ghosted”) pending crew time availability.]

Kaleri’s disassembly of the VELO cycle ergometer yesterday did not locate the cause of the previous failure of its control module nor did it uncover the reason of its sudden recovery after a “kick”.  The unit was reassembled OK and is continuing to function well.

MCC-H EVA specialists reported that all three on-orbit EMU/spacesuits (## 3005, 3011, 3013) have been cleared of O2 contamination concerns. [##3005 and 3011 are Go for EVA operations, while EMU #3013 remains without cooling flow and thus is No-Go.]

TsUP/Moscow today completed a one-day test on the FGB solar arrays, switching them back to primary power channel after taking them to the secondary channel yesterday.

TVIS Update:  The crew today inspected the TVIS treadmill to assess the wire ropes in its four corners, taking multiple digital imagery.  Condition of the wire ropes was found unchanged from the previous inspection (which discovered two ropes heavily frayed), i.e., still “within Flight Rules”, so the treadmill continues to be Go for exercise without detriment to crew workout performance.  Failure of the stabilizing corner wire ropes, used for centering the frame, would not be a safety issue but would impact vibration isolation.  The video imagery from today’s TVIS testing is to be downlinked tomorrow for review; afterwards, further steps concerning the ropes will be discussed early next week.  Then, IFM (inflight maintenance) activities on TVIS will be conducted on two days, with 4 hours each day: On 12/9 (Tuesday), the crew will remove TVIS from the SM floor pit for a close look at the failed roll-axis stabilizing gyro, to identify the failure so that repair tools and replacement parts can be specified for Progress 13P.  They will also reconfigure the yaw and pitch stabilizers, by resetting “dipswitches”, to allow their function during exercising with the gyro unpowered.  Results of the IFM will be studied on the ground overnight, followed on 12/10 by TVIS reinstallation in the pit and closeout ops.  On 12/16, time will be budgeted in case TVIS requires additional work.

Structural dynamics (vibration) data are being taken during TVIS testing with the IWIS (internal wireless instrumentation system), with all three RSUs (remote sensing units) active, two in the SM, one in the Node.  The latter, RSU #1027, found nonfunctional on 11/24, was recovered and is back in operation. [RSUs are radio-linked accelerometers in SM, Lab and Node that take structural dynamics measurements and transmit them to a central control station in the Lab.]

CMG Update:  Status of and plans for the three U.S. control moment gyros (CMGs) were discussed with Russia yesterday at a joint operations planning (JOP) meeting.  Both sides agreed that ISS attitude hold and all attitude maneuvers will be performed on Russian thrusters only between now and mid-January.  The CMGs will be available for attitude hold if needed, but their use for this will be kept at a minimum.  Meanwhile, specialist teams will evaluate CMG data and provide new recommendations for CMG constraints by mid-January.  Propellant use for this ops mode still needs to be assessed by Moscow but is anticipated to be small (“less than 25 kg”), a minor impact in light of the large propellant amounts currently stored in the FGB, SM and Progress (~ 3700 kg).

Consumables Update:  Onboard consumables remain in good shape, generally tracking predictions quite well.  Water use remains slightly above the 2 liters/day projection, and four EDVs with H2O will be manifested on Progress 13P.  Food, both U.S. and Russian, will also have to be shipped up on 13P.  Propellant supplies are more than adequate (see above), particularly since an originally planned reboost in October was not performed.  13P manifesting is also being considered for new RED (resistive exercise device) canisters as well as for the to-be-determined TVIS hardware (see above).  Procedures to be taken for best utilization of the O2 (oxygen) supply stored in Progress 12P are being discussed with Moscow.

Strange Noise Update:  Additional external viewing activities with the video cameras of the SSRMS (Space Station Remote Manipulator System) are being scheduled for 12/11 and 12/18.  Surveys will probably include the FGB, SM portside and zenith (top) surfaces, P6 truss 4-bar port shoulder bolt, and ISS top side.

Today’s CEO targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, were Madras (Chennai), India (pointing just left for this coastal city that is becoming the hub for India’s computer industry), Calcutta, India(pointing left on the west margin of the Ganges delta.  No good handheld images exist of this major metropolis of 13.39 million), Ganges River Delta (physical change of complex coastlines is the science focus of this site.  But land use, especially agricultural and logging incursions into the protected forests along the coast, are of high ecological interest), Dhaka, Bangladesh (nadir pass over this capital city of 13.43 million), Cape Town, South Africa (nadir pass),Karachi, Pakistan (nadir pass over Pakistan’s largest city, of 12.64 million), Lima, Peru (nadir pass), Mexico City, Mexico (pointing a touch left), and Albuquerque, New Mexico (pointing a touch right).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status (as of 2:00pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On, 20A.  Vozdukh CO2 scrubber is On (in Manual Mode 5/3).  U.S. CDRA CO2 scrubber is on Standby (ready in dual-bed mode).  TCCS (trace contaminant control subsystem) is operating.  MCA (major constituents analyzer) is off (in Life Extending Mode).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On; SKV-2 is Off.
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 26.5; ppO2 (mmHg) — 158.5; ppCO2 (mmHg) — 3.9.
  • SM Transfer Compartment:  Pressure (mmHg) — 761; temperature (deg C) — 20.2.
  • FGB Cabin:  Pressure (mmHg) — 756; temperature (deg C) — 22.3.
  • Node:  Pressure (mmHg) — 754.43; temperature (deg C) — 21.9 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 755.62; temperature (deg C) — 24.9; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock): Pressure (mmHg) — 755.72; temperature (deg C) — 22.9; shell heater temp (deg C) — 23.3, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1:  Shell heater temp (deg C) — 21.9
  • PMA-2:  Shell heater temp (deg C) — 19.0

(n/a = data not available)

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in fixed (dual-angle) position (non-suntracking). 
  • SM batteries:  Battery #8, formerly known as #7, is still in slot #8 for troubleshooting; all other batteries (7) are in “Partial Charge” mode.
  • FGB batteries:  Battery #5 is off; all other batteries (5) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Command & Data Handling Systems:

  • C&C-2 MDM is prime, C&C-1 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational (all lanes reintegrated 11/5).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational; string #3 dropped out 10/22.

Propulsion System:

  • Total propellant load available:3704 kg (8166 lb) as of 11/13 [SM(755) + FGB(2597) + Progress M(352) + Progress M-1(0)].  (Capability: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH  XVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, x-axis in velocity vector [yaw: -10 deg, pitch: -9.4 deg, roll: 0 deg]), with CMG/TA (thruster assist) Momentum Management.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-2 is prime, IAC-1 is suspect).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.

Robotics:

  • SSRMS/Canadarm2 based at MBS PDGF #2/LEE A, with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 7:39am EST [= epoch]):

  • Mean altitude — 372.5 km
  • Apogee — 376.7 km
  • Perigee — 368.4 km
  • Period — 92.0 min.
  • Inclination (to Equator) —  51.63 deg
  • Eccentricity — 0.000611
  • Orbits per 24-hr. day — 15.65
  • Mean altitude loss in last 24 hours — 110 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 28763
  • For more on ISS orbit and worldwide naked-eye visibility dates/times, see
  •  http://www.hq.nasa.gov/osf/station/viewing/issvis.html

SpaceRef staff editor.